# Artin Symmetric Functions

### v2

@@ 1 Introduction
In this paper, we will introduce a family of invariants attached to Galois representations called the Artin symmetric functions $\operatorname{\mathscr{L}}(\operatorname{\mathfrak{X}}_{K},\rho;L/K).$ These are algebraic-combinatorial objects which live in the ring of arithmetic symmetric functions associated to the base field $K.$ This ring is a graded completion $\mathcal{A}_{K}$ of a ring, which we write as $A_{K}$, defined as an infinite tensor product of rings of symmetric functions in variables $\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}$ labelled by prime ideals $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}$ of the ring of integers $\mathcal{O}_{K}$. On its face, the rings $A_{K}\subset\mathcal{A}_{K}$ are not very interesting. However, as is common in algebraic combinatorics with the study of ordinary symmetric functions $\Lambda$, we find interesting structure inside the rings $A_{K}\subset\mathcal{A}_{K}$ by looking at special bases defined combinatorially. Specifically, the ring $A_{K}$, has a basis $\{P_{\lambda^{\bullet}}[\operatorname{\mathfrak{X}}_{K}]\}_{\lambda^{\bullet}\in\mathbb{Y}_{K}}$ indexed by almost everywhere empty sequences of partitions $\lambda^{\bullet}=(\lambda^{(\operatorname{\mathfrak{q}})})_{\operatorname{\mathfrak{q}}\in\mathbb{Y}_{K}}$ called the arithmetic Hall-Littlewood functions which are defined as products of ordinary Hall-Littlewood $P$-functions $P_{\lambda^{(\operatorname{\mathfrak{q}})}}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})};N(\operatorname{\mathfrak{q}})^{-1}].$ This construction is highly inspired by Chapter IV of Macdonald’s book [Macdonald] where a similar collection of functions is constructed in the setting of $\mathbb{F}_{q}((t))$ to describe the characters of $\operatorname{\mathrm{GL}}_{n}(\mathbb{F}_{q}).$ We develop a basic theory for $\mathcal{A}_{K}$ with respect to the basis $P_{\lambda^{\bullet}}[\operatorname{\mathfrak{X}}_{K}]$ and relate this theory to the finite global spherical Hecke algebras $\mathcal{H}_{n}(K)$ for the locally compact groups $\operatorname{\mathrm{GL}}_{n}(\widehat{\mathcal{O}}_{K})\backslash\operatorname{\mathrm{GL}}_{n}(\mathbb{A}_{K}^{f})/\operatorname{\mathrm{GL}}_{n}(\widehat{\mathcal{O}}_{K})$ for $n\geq 1.$
@@
1.1 Acknowledgements
The author would like to thank Daniel Orr and Monica Vazirani for helpful discussions about Hall-Littlewood polynomials and Hecke algebras.
@@ 2.2 Arithmetic Symmetric Functions
For the majority of this paper we will work over $\mathbb{Q}^{\text{ab}}$ except for during the last section when we will be proving analytic results. The choice of $\mathbb{Q}^{\text{ab}}$ over, the possibly more natural choice of, $\mathbb{C}$ is to emphasize that the arithmetic symmetric functions are not analytic objects (on their own) and so we may take coefficients over any commutative ring. Most relevant for studying Galois representations, one could consider arithmetic symmetric functions over an $\ell$-adic field $E/\mathbb{Q}_{\ell}$ and talk about $\ell$-adic Galois representations. We will not do so in this paper however.
@@
For $\mathfrak{a}\in\operatorname{\mathcal{J}}_{K}$ define $\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a})\in\mathbb{Z}$ using unique factorization by
@@
$H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]:=h_{m}\left[\sum_{\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}\right]\in A_{K}.$
@@
$\operatorname{\mathrm{Exp}}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]:=\operatorname{\mathrm{Exp}}\left[\sum_{\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}\right]=\sum_{m\geq 0}H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]\in\mathcal{A}_{K}.$
Note that the $H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]$ are not always homogeneous with respect to the grading ||||.|$\cdot$||.| | ⋅ | | . However, if we define the simpler grading $\deg^{\prime}$ on $A_{K}$ given by the degree in the $x_{i}^{(\operatorname{\mathfrak{q}})}$ variables (without distinguishing between distinct $\operatorname{\mathfrak{q}}$), then the $H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]$ are homogeneous with $\deg^{\prime}(H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}])=m.$ The following properties are straightforward to check.
@@
  • The ring $\mathcal{O}_{K}$ is a unique factorization domain if and only if $\mathscr{CL}_{K}=\mathbb{Q}^{\text{ab}}$.
@@
For the first statement, suppose that $\mathcal{O}_{K}$ is a UFD. Then for each integral prime ideal $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}$ and $m\geq 1$, $h_{m}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]=H^{(m)}_{\operatorname{\mathfrak{q}}}\in\mathscr{P}_{K}.$ The set $\{h_{m}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]\leavevmode\nobreak\ |\leavevmode\nobreak\ m\geq 1,\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}\}$ generates $A_{K}$ algebraically so $\mathscr{CL}_{K}=A_{K}/\mathscr{P}_{K}=\mathbb{Q}^{\text{ab}}.$ Conversely, if $\mathscr{CL}_{K}=\mathbb{Q}^{\text{ab}}$ then by the algebraic independence of the homogeneous symmetric functions $h_{m}[\operatorname{\mathfrak{X}}]\in\Lambda$ for $m\geq 1$, we know that $H^{(m)}_{\operatorname{\mathfrak{q}}}=h_{m}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]\in\mathscr{P}_{K}$ for all $m\geq 1$ and $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}.$ Therefore, by the definition of $\mathscr{P}_{K}$ we must have in particular that $h_{1}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]$ is an $A_{K}$ linear combination of elements of the form $H^{(m)}_{(\alpha)}$ for $m\geq 1$ and $\alpha\in K^{*}.$ By considering the degree of each $\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{p}})}$ variable set separately the only way this is possible is if $h_{1}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]=H^{(1)}_{(\alpha)}$ for some $\alpha\in K^{*}$ which means that $\operatorname{\mathfrak{q}}=(\alpha).$ Thus $\mathcal{O}_{K}$ is a UFD.
@@
It is an important result in classical algebraic number theory that the ideal class group $\mathrm{Cl}_{K}$ of any number field K is a finite group. We may show an analogue of this result for the ring $\mathscr{CL}_{K}$. Note that the rings $A_{K},\mathcal{A}_{K}$ are $\lambda$-rings with the ordinary plethysm operators $p_{r}$ acting in the variables as $x_{i}^{(\operatorname{\mathfrak{q}})}\rightarrow(x_{i}^{(\operatorname{\mathfrak{q}})})^{r}$ and extended accordingly.
@@
First, we argue that the $\lambda$-ring structure on $A_{K}$ descends to $\mathscr{CL}_{K}.$ Since over the field $\mathbb{Q}\subset\mathbb{Q}^{\text{ab}}$ the power sum symmetric functions $p_{r}$ for $r\geq 1$ generate $\Lambda$ algebraically,
$\mathscr{P}_{K}=\langle H_{(\alpha)}^{(m)}\leavevmode\nobreak\ |\leavevmode\nobreak\ m\geq 1,\alpha\in K^{*}\rangle=\langle P_{(\alpha)}^{(r)}\leavevmode\nobreak\ |\leavevmode\nobreak\ r\geq 1,\alpha\in K^{*}\rangle$
where
$P_{\mathfrak{a}}^{(r)}:=p_{r}\left[\sum_{\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}\right].$
Let $r\geq 1$ and take $F=\sum_{i}g_{i}P_{(\alpha_{i})}^{(m_{i})}\in\mathscr{P}_{K}$ for $\alpha_{i}\in K^{*}.$ Then
$p_{r}[F]=p_{r}\left[\sum_{i}g_{i}P_{(\alpha_{i})}^{(m_{i})}\right]=\sum_{i}p_{r}[g_{i}]p_{r}[P_{(\alpha_{i})}^{(m_{i})}]=\sum_{i}p_{r}[g_{i}]p_{rm_{i}}\left[\sum_{\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}\right]=\sum_{i}p_{r}[g_{i}]P_{(\alpha_{i})}^{(rm_{i})}\in\mathscr{P}_{K}.$
Therefore, $\mathscr{P}_{K}$ is a $\lambda$-ideal of $A_{K}$ and so $\mathscr{CL}_{K}$ is a $\lambda$-ring. Next we appeal to the classical result that $\mathrm{Cl}_{K}$ is a finite group. Let $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{N}$ be a finite set of representatives for the distinct (fractional) ideal classes in $\mathrm{Cl}_{K}.$ We claim that the elements $X_{i}:=\sum_{\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a}_{i})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}$ generate $\mathscr{CL}_{K}$ as a $\lambda$-ring. Since $A_{K}$ is generated by the elements $p_{r}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]$ for $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}$, it suffices to show that $p_{r}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]$ is contained in the $\lambda$-ring generated by $X_{1},\ldots,X_{N}.$ Let $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}$. By assumption $\operatorname{\mathfrak{q}}=(\alpha)\mathfrak{a}_{i}$ for some $1\leq i\leq N$ and $\alpha\in K^{*}.$ Then
@@
$\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}=\sum_{\operatorname{\mathfrak{p}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{p}}}(\mathfrak{a}_{i})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{p}})}=X_{i}.$
But now we see that $p_{r}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]=p_{r}[X_{i}]$ in $\mathscr{CL}_{K}$ and so $p_{r}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]$ is contained in the $\lambda$-ring generated by $X_{1},\ldots,X_{r}$. Therefore, $\mathscr{CL}_{K}$ is finitely generated as a $\lambda$-ring. ∎

### v3

@@ 1 Introduction
In this paper, we will introduce a family of invariants attached to Galois representations called the Artin symmetric functions $\operatorname{\mathscr{L}}(\operatorname{\mathfrak{X}}_{K},\rho;L/K).$ These are algebraic-combinatorial objects which live in the ring of arithmetic symmetric functions associated to the base field $K.$ This ring is a graded completion $\mathcal{A}_{K}$ of a ring, which we write as $A_{K}$, defined as an infinite tensor product of rings of symmetric functions in variables $\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}$ labelled by prime ideals $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}$ of the ring of integers $\mathcal{O}_{K}$. On its face, the rings $A_{K}\subset\mathcal{A}_{K}$ are not very interesting. However, as is common in algebraic combinatorics with the study of ordinary symmetric functions $\Lambda$, we find interesting structure inside the rings $A_{K}\subset\mathcal{A}_{K}$ by looking at special bases defined combinatorially. Specifically, the ring $A_{K}$, has a basis $\{P_{\lambda^{\bullet}}[\operatorname{\mathfrak{X}}_{K}]\}_{\lambda^{\bullet}\in\mathbb{Y}_{K}}$ indexed by almost everywhere empty sequences of partitions $\lambda^{\bullet}=(\lambda^{(\operatorname{\mathfrak{q}})})_{\operatorname{\mathfrak{q}}\in\mathbb{Y}_{K}}$ called the arithmetic Hall-Littlewood functions which are defined as products of ordinary Hall-Littlewood $P$-functions $P_{\lambda^{(\operatorname{\mathfrak{q}})}}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})};N(\operatorname{\mathfrak{q}})^{-1}].$ This construction is highly inspired by Chapter IV of Macdonald’s book [Macdonald] where a similar collection of functions is constructed in the setting of $\mathbb{F}_{q}(t)$ to describe the characters of $\operatorname{\mathrm{GL}}_{n}(\mathbb{F}_{q}).$ We develop a basic theory for $\mathcal{A}_{K}$ with respect to the basis $P_{\lambda^{\bullet}}[\operatorname{\mathfrak{X}}_{K}]$ and relate this theory to the finite global spherical Hecke algebras $\mathcal{H}_{n}(K)$ for the locally compact groups $\operatorname{\mathrm{GL}}_{n}(\widehat{\mathcal{O}}_{K})\backslash\operatorname{\mathrm{GL}}_{n}(\mathbb{A}_{K}^{f})/\operatorname{\mathrm{GL}}_{n}(\widehat{\mathcal{O}}_{K})$ for $n\geq 1.$
@@
1.1 Acknowledgments
The author would like to thank Daniel Orr and Monica Vazirani for helpful discussions about Hall-Littlewood polynomials and Hecke algebras. The author would also like to thank Sebastian Bozlee and Leo Herr for interesting conversations about Picard groups and for pointing out an error in earlier versions of Propositions 2.23 2.24 as well as thank Gergely Harcos for a correction to the introduction of a previous version of this paper.
@@ 2.2 Arithmetic Symmetric Functions
For the majority of this paper we will work over $\mathbb{Q}^{\text{ab}}$ except for a few sections and we will make the change clear. The choice of $\mathbb{Q}^{\text{ab}}$ over, the possibly more natural choice of, $\mathbb{C}$ is to emphasize that the arithmetic symmetric functions are not analytic objects (on their own) and so we may take coefficients over any commutative ring. Most relevant for studying Galois representations, one could consider arithmetic symmetric functions over an $\ell$-adic field $E/\mathbb{Q}_{\ell}$ and talk about $\ell$-adic Galois representations. We will not do so in this paper however.
@@
Write $A_{K}(\mathbb{Z})$ and $\mathcal{A}_{K}(\mathbb{Z})$ for the rings defined as in Definitions 2.12 2.14 analogously but with coefficients in $\mathbb{Z}$ instead of $\mathbb{Q}^{\text{ab}}.$ We will similarly write $\Lambda_{\mathbb{Z}}$ for the ring of symmetric functions over $\mathbb{Z}.$ For $\mathfrak{a}\in\operatorname{\mathcal{J}}_{K}$ define $\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a})\in\mathbb{Z}$ using unique factorization by
@@
$H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]:=h_{m}\left[\sum_{\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}\right]\in A_{K}(\mathbb{Z}).$
@@
$\operatorname{\mathrm{Exp}}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]:=\operatorname{\mathrm{Exp}}\left[\sum_{\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}\right]=\sum_{m\geq 0}H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]\in\mathcal{A}_{K}(\mathbb{Z}).$
Note that the $H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]$ are not always homogeneous with respect to the grading ||||.|$\cdot$||.| | ⋅ | | . However, if we define the simpler grading $\deg^{\prime}$ on $A_{K}(\mathbb{Z})$ given by the degree in the $x_{i}^{(\operatorname{\mathfrak{q}})}$ variables (without distinguishing between distinct $\operatorname{\mathfrak{q}}$), then the $H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}]$ are homogeneous with $\deg^{\prime}(H^{(m)}_{\mathfrak{a}}[\operatorname{\mathfrak{X}}_{K}])=m.$ The following properties are straightforward to check.
@@
  • The ring $\mathcal{O}_{K}$ is a unique factorization domain if and only if $\mathscr{CL}_{K}=\mathbb{Z}$.
@@
For the first statement, suppose that $\mathcal{O}_{K}$ is a UFD. Then for each integral prime ideal $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}$ and $m\geq 1$, $h_{m}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]=H^{(m)}_{\operatorname{\mathfrak{q}}}\in\mathscr{P}_{K}.$ The set $\{h_{m}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]\leavevmode\nobreak\ |\leavevmode\nobreak\ m\geq 1,\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}\}$ generates $A_{K}(\mathbb{Z})$ algebraically so $\mathscr{CL}_{K}=A_{K}(\mathbb{Z})/\mathscr{P}_{K}=\mathbb{Z}.$ Conversely, if $\mathscr{CL}_{K}=\mathbb{Z}$ then by the algebraic independence of the homogeneous symmetric functions $h_{m}[\operatorname{\mathfrak{X}}]\in\Lambda$ for $m\geq 1$, we know that $H^{(m)}_{\operatorname{\mathfrak{q}}}=h_{m}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]\in\mathscr{P}_{K}$ for all $m\geq 1$ and $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}.$ Therefore, by the definition of $\mathscr{P}_{K}$ we must have in particular that $h_{1}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]$ is an $A_{K}(\mathbb{Z})$ linear combination of elements of the form $H^{(m)}_{(\alpha)}$ for $m\geq 1$ and $\alpha\in K^{*}.$ By considering the degree of each $\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{p}})}$ variable set separately the only way this is possible is if $h_{1}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]=H^{(1)}_{(\alpha)}$ for some $\alpha\in K^{*}$ which means that $\operatorname{\mathfrak{q}}=(\alpha).$ Thus $\mathcal{O}_{K}$ is a UFD.
@@
It is an important result in classical algebraic number theory that the ideal class group $\mathrm{Cl}_{K}$ of any number field K is a finite group. We may show an analogue of this result for the ring $\mathscr{CL}_{K}$. Note that the rings $A_{K}(\mathbb{Z}),\mathcal{A}_{K}(\mathbb{Z})$ are $\lambda$-rings with the ordinary plethysm operators $h_{r}$ and extended accordingly. Equivalently, we may consider the power sum plethysm operators $p_{r}:x_{i}^{(\mathfrak{q})}\rightarrow(x_{i}^{(\mathfrak{q})})^{r}$ for all $r\geq 1$ and expand each $h_{r}$ into the $p_{\lambda}$ accordingly. However, these expansions require coefficients in $\mathbb{Q}$ (for example $2h_{2}=p_{1,1}+p_{2}$) but luckily the effect of computing $h_{r}[F]$ for any $F\in\mathcal{A}_{K}(\mathbb{Z})$ will still have integral coefficients.
@@
First, we argue that the $\lambda$-ring structure on $A_{K}(\mathbb{Z})$ descends to $\mathscr{CL}_{K}.$ Note that over $\mathbb{Z},$ the complete homogeneous symmetric functions $h_{r}$ for $r\geq 1$ generate $\Lambda_{\mathbb{Z}}$ so it suffices to show that for $r\geq 1$ and any arbitrary element $F=\sum_{i=1}^{k}g_{i}H_{(\alpha_{i})}^{(m_{i})}\in\mathscr{P}_{K}$ we have that $h_{r}[F]\in\mathscr{P}_{K}.$ To this end we have that
$h_{r}[F]=h_{r}\left[\sum_{i=1}^{k}g_{i}H_{(\alpha_{i})}^{(m_{i})}\right]=\sum_{a_{1}+\ldots+a_{k}=r}\prod_{i=1}^{k}h_{a_{i}}\left[g_{i}H_{(\alpha_{i})}^{(m_{i})}\right]=\sum_{a_{1}+\ldots+a_{k}=r}\prod_{i=1}^{k}\left(\sum_{|\lambda|=a_{i}}s_{\lambda}[g_{i}]s_{\lambda}\left[H_{(\alpha_{i})}^{(m_{i})}\right]\right).$
By the Jacobi-Trudi identity, each $s_{\lambda}$ expands as a $\mathbb{Z}$-linear combination of the $h_{\mu}$ functions. Thus it suffices to show that for all $\ell\geq 1$ and $1\leq i\leq k$, $h_{\ell}\left[H_{(\alpha_{i})}^{(m_{i})}\right]\in\mathscr{P}_{K}.$ We see that
$h_{\ell}\left[H_{(\alpha_{i})}^{(m_{i})}\right]=h_{\ell}\left[h_{m_{i}}\left[\sum_{\mathfrak{q}\in\mathbb{P}_{K}}\nu_{\mathfrak{q}}(\alpha_{i})\operatorname{\mathfrak{X}}^{(\mathfrak{q})}\right]\right]=\left(h_{\ell}[h_{m_{i}}]\right)\left[\sum_{\mathfrak{q}\in\mathbb{P}_{K}}\nu_{\mathfrak{q}}(\alpha_{i})\operatorname{\mathfrak{X}}^{(\mathfrak{q})}\right]$
which now certainly lives in the ideal $\mathscr{P}_{K}$. Therefore, $\mathscr{P}_{K}$ is a $\lambda$-ideal of $A_{K}(\mathbb{Z})$ and so $\mathscr{CL}_{K}$ is a $\lambda$-ring.
Next we appeal to the classical result that $\mathrm{Cl}_{K}$ is a finite group. Let $\mathfrak{a}_{1},\ldots,\mathfrak{a}_{N}$ be a finite set of representatives for the distinct (fractional) ideal classes in $\mathrm{Cl}_{K}.$ We claim that the elements $B_{i}:=\sum_{\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{q}}}(\mathfrak{a}_{i})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}$ generate $\mathscr{CL}_{K}$ as a $\lambda$-ring. Since $A_{K}(\mathbb{Z})$ is generated by the elements $h_{r}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]$ for $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}$, it suffices to show that $h_{r}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]$ is contained in the $\lambda$-ring generated by $B_{1},\ldots,B_{N}.$ Let $\operatorname{\mathfrak{q}}\in\mathbb{P}_{K}$. By assumption $\operatorname{\mathfrak{q}}=(\alpha)\mathfrak{a}_{i}$ for some $1\leq i\leq N$ and $\alpha\in K^{*}.$ Then
@@
$\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}=\sum_{\operatorname{\mathfrak{p}}\in\mathbb{P}_{K}}\nu_{\operatorname{\mathfrak{p}}}(\mathfrak{a}_{i})\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{p}})}=B_{i}.$
But now we see that $h_{r}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]=h_{r}[B_{i}]$ in $\mathscr{CL}_{K}$ and so $h_{r}[\operatorname{\mathfrak{X}}^{(\operatorname{\mathfrak{q}})}]$ is contained in the $\lambda$-ring generated by $B_{1},\ldots,B_{r}$. Therefore, $\mathscr{CL}_{K}$ is finitely generated as a $\lambda$-ring. ∎